raju is now 4 times as old as sony.five years ago raju was 7 times as old as sony.find their present age.
Answer by reviewermath(1029)

Let x = current age of Sony
so 4x = current age of Raju
x - 5 = age of Sony 5 years ago
4x - 5 = age of Raju 5 years ago
4x - 5 = 7(x - 5)
4x - 5 = 7x - 35
-3x = -30
x = 10
Therefore, Sony is 10 years old and Raju is 40 years old.